SW&IT Meeting Minutes 2019-01-07
Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang
Report auto-generated every week, Do not edit this table, use entries in smartsheets.
| LR20181217 - Bec, Matthieu
| LR20181217 - Colson, Andrew
| LR20181217 - Edmeades, David
| LR20181217 - Sola, Igor
| LR20181217 - Kubanek, Petr
| LR20181217 - Hooper, Stephen
| LR20181217 - Zhang, Yang
| LR20181217 - Zhang, Xianyu
| LR20181231 - Bec, Matthieu
| LR20181231 - Colson, Andrew
| LR20181231 - Edmeades, David
| LR20181231 - Sola, Igor
| LR20181231 - Kubanek, Petr
| LR20181231 - Hooper, Stephen
| LR20181231 - Zhang, Yang
| LR20181231 - Zhang, Xianyu
News from management
- PEPSI/pol high profile run in January: AGW3 & 4 must be ready
Previous Action Items
Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.
- network dropout - antenna monitoring?
- iLocater Single-Mode Data Fibers - CRA/CRB to Electronics Shop
- BCU f/w upgrade documentation
- revert simbad harvard-strasbourg /et/hosts hack from Friday
- Dust monitor - bring back
- AO improvement: Luci kwds
- AO cabinet switch upgrade - follow up
- wifi summit upgrade
- Tucson CR power overhaul plan
- n507/obs upgrade
- abell / jhill ws
- synology new drives
- PGray mirror cover req. feedback
- NX on obs3
- OAC config, version
- VMware alarm emails
- clone of weather station winxp?
- Mangement, OC, WPC
- engineering science archive (ESA) prototyping on analysis, making good use of `parallel`
- misc: inbox clean up, macos Mojave upgrade
- phone screening: Jake, Anil - 3rd has not responded
- issues with Smartsheet and API: timeout, errors. contacted support
- wiki Software/WebHome refactoring
- OAC / PEPSI support
- git migrations
- Jake Weiss zoom interview Thu @ 9am
- new candidates screening
- ESA wip continued
- finish up alarm handling system refactoring (AHS in git)
- LBC github migration.
- IT issue #7517 Initializing wfs cam process stuck on GCSL (MODS1) => monitoring.
- IT issue #7518 NaN pixels in LBCR images, in discussion with Olga to determine whether to change LBC image fits header info.
- Put fits file to PNG thumbnail python script to Github.
- Working on AO-Cravat server part (setting up mod_python).
- Moved 2018 telemetry to safekeeping dir, started new "tcs" telemetry.
- Bucki's laptop accepted, installed, delivered to client.
- New 507 Obs workstations delivered in Kellee's office.
- Test of second port on Moxa.
- Email followups from break
- Notably, everyone seems to be good on Autodesk software licenses
- I will follow up special with Jonathan to make sure that he's on the same page as everyone else
- Start work on power plan using existing hardware as much as possible
- VMWare host warning investigation
- AZCam and weatherstation is now in Git (MB)
- Particle monitor telemetry
- ARGOS headers
- Added loop frequency, mode, and radians
- need confirmation of frequency
- pending PR
- mountain configuration file updated (hostname, HK port and IP, and hostname)
- Consider JSON for ascii format output of getdata
- Release 5.1 to the mountain
- ARGOS on-call support 1/8
- ESM/ARGOS converston script
- Subwindow debugging: Trying to understand a bit better GEIRS backend interaction in SRR mode also part of #7473
- Looking at #7425 Filter wheel issue
- OVMS+TT meeting and wrote the SW ICD.
- 4D interferometer HW requirement
- EU visa paper work